OpenStudy (anonymous):

Ammon and Nika volunteer at an animal shelter, Nakia worked 3 more hours than Ammon. They each worked a whole number of hours. They together worked more than 27 hours. What is the least number of hours each worked?

OpenStudy (dan815):

both of they hours added is 27 and one of them worked 3 hrs more

jimthompson5910 (jim_thompson5910):

x = # of hours Ammon worked y = # of hours Nika worked "Nakia worked 3 more hours than Ammon" so y = x + 3 is your first equation

OpenStudy (dan815):

their*

jimthompson5910 (jim_thompson5910):

" They together worked more than 27 hours" so the sum of their hours, x+y, has to be more than 27 meaning that x+y > 27

jimthompson5910 (jim_thompson5910):

x+y > 27 x+x+3 > 27 ... replace y with x+3 (works because y = x+3) solve for x

OpenStudy (anonymous):

\[is \it 5x >27\]

OpenStudy (anonymous):

is it*

jimthompson5910 (jim_thompson5910):

should be 2x+3 > 27 2x > 27-3 2x > 24 x > ???

OpenStudy (anonymous):

12?

jimthompson5910 (jim_thompson5910):

yep x > 12

OpenStudy (anonymous):

ok ty

jimthompson5910 (jim_thompson5910):

so recall that we said that "x = # of hours Ammon worked" so x > 12 means Ammon worked more than 12 hrs the min Ammon can work is therefore 13 hours (he worked a whole number of hours)

jimthompson5910 (jim_thompson5910):

use x = 13 to find y

OpenStudy (anonymous):

how

jimthompson5910 (jim_thompson5910):

y = x+3 is the first equation we set up

jimthompson5910 (jim_thompson5910):

plug in x = 13

OpenStudy (anonymous):

thanks i got the answer

OpenStudy (anonymous):

I have another problem

jimthompson5910 (jim_thompson5910):

alright, I'll help with one more

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

its a word problem

OpenStudy (anonymous):

The length of a rectangle is 4cm longer than the width, and the perimeter is at least 48 cm. What are the smallest possible dimensions of the rectangle?

jimthompson5910 (jim_thompson5910):

"The length of a rectangle is 4cm longer than the width" L = W+4 L is the length, W is the width

jimthompson5910 (jim_thompson5910):

"perimeter is at least 48 cm" P >= 48 2L+2W >= 48 2(W+4)+2W >= 48 solve for W

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

\[w \ge10\]

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

now plug w = 10 into L = w+4

OpenStudy (anonymous):

L=14

jimthompson5910 (jim_thompson5910):

L = 14 is correct